OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form.

Directrix (directrix):

Point-slope form of a line y - y1 = m* (x - x1) where m is the slope of the line and (x1,y1) is the point (x1,y1) = (–1, 5) m = 3 Fill in for (x1,y1) and m in y - y1 = m* (x - x1) @Tamar99

OpenStudy (anonymous):

thank you @Directrix

Directrix (directrix):

You are welcome. What did you get for the answer? @Tamar99